Chiwen, Inc. is considering whether to replace an office copy machine with a new model.  The manager is trying to decide whether to buy the machine outright or lease it from a copier company.  He has generated the following information about the two options.

Purchase Option

Lease Option

Purchase price

$ 25,000

Annual lease payment (includes all supplies and maintenance)

$ 7,000

Annual operating costs (paper, toner, maintenance, etc.)

$ 2,000

Useful life

5 years

Contract length

5 years

Salvage value

$ 5,000

Chiwen’s discount rate is 12 percent.  Should the company lease or buy the copier?
Answer #1

(i) To purchase of new machine

Amount ($)

Present value of after-tax cash outflows

Cost of purchase

Equivalent annual cost for 5 years = 25000

3.605

6934.81

Running and maintenance cost per annum

5000.00

Total net equivalent cash outflows p.a.

11934.81

(ii) Lease a new machine

Amount ($)

Lease Rent

7000.00

Total net equivalent cash outflows p.a.

7000.00

(iii) Since, net equivalent cash outflows p.a. for buying a new machine $ 11934.81 is HIGHER than net equivalent cash outflows of $ 7000.00 for Lease a new machine. Therefore, it is advisable that the company should go for Lease a new machine.

WORKING NOTES

1PVAF5/.12=1(1(1+.12)−5.12)=1(10.5674268557186.12)=1(3.604776202345)=3.604776202345